40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2883     Entry time: Wed May 5 16:58:21 2010
Author: Koji 
Type: Update 
Category: PSL 
Subject: 2W hooked up to the interlock service 

Ben, Steve, and Koji

Ben came to the 40m and hooked up a cable to the main interlock service.
We have tested the interlock and confirmed it's working.

[Now the laser is approved to be used by persons who signed in the SOP.]

The RC, PMC, and MZ were unlocked during the interlock maneuver.
Now they are relocked.

ELOG V3.1.3-